USING TRYPTOPHAN AS A PROBE FOR STUDYING PROTEIN HYDRATION DYNAMICS

\underline{YANGZHONG QIN}, CHIH-WEI CHANG, LIJUAN WANG AND DONGPING ZHONG, DEPART-MENT OF PHYSICS, THE OHIO STATE UNIVERSITY, COLUMBUS, OH, 43210.

Although tryptophan has been used as a probe to study solvation dynamics for decades, a few questions still remain unclear, such as missing of the initial fast component. We proposed a solvation model for construction of the response function of tryptophan probe and carefully investigated the possible initial loss of the ultrafast component with established time-zero emission spectrum and femtosecond-resolved emission spectra (FRES)^{a}. With mutation of GB1 protein, we demonstrated that tryptophan is a powerful probe for protein solvation dynamics and show significant slowdown of the dynamics of hydration shell at protein-water interface.

aY. Qin, C.-W. Chang, L. Wang and D. Zhong, J. Phys. Chem. B. \underline{116}, 13320, (2012).